Suppose that a population yy grows according to the logistic model given by formula:



yy = LL


1 + AAee−kkkk .


a. At what rate is yy increasing at time tt = 0 ?


b. In words, describe how the rate of growth of yy varies with time.


c. At what time is the population growing most rapidly?

1
Expert's answer
2021-11-29T14:59:25-0500

"y=\frac{l}{1+Ae^{-kt}}"


a) "\frac{dy}{dt}=\frac{-l\cdot Ae^{-kt}(-k)}{(1+Ae^{-kt})^2}"

"=\frac{klAe^{-kt}}{(1+Ae^{-kt})^2}"

"y'(t=0)=\frac{kAl}{(1+A)^2}"

y is increasing at a rate of "\frac{kAl}{(1+A)^2}"

b) The rate of growth is always positive meaning y will increase with increase in t

c) "y''=\frac{(1+Ae^{-kt})^2(-k^2Ale^{-kt})-kAle^{-kt}\cdot 2(1+Ae^{-kt})(-kA)}{(1+Ae^{-kt})^4}=0"

"\implies(1+Ae^{-kt})(-1)+(+2)=0"

"\implies 1+Ae^{-kt}=2"

"e^{-kt}=\frac{1}{A}"

"t=\frac{ln{\frac{1}{A}}}{-k}"

Population is growing most rapidly at "t=\frac{ln{\frac{1}{A}}}{-k}"